OpenStudy (anonymous):

A sequence has its first term equal to 3, and each term of the sequence is obtained by adding 5 to the previous term. If f(n) represents the nth term of the sequence, which of the following recursive functions best defines this sequence?

OpenStudy (anonymous):

f(1) = 3 and f(n) = f(n - 1) + 5; n > 1 f(1) = 5 and f(n) = f(n - 1) + 3; n > 1 f(1) = 3 and f(n) = f(n - 1) + 5n; n > 1 f(1) = 5 and f(n) = f(n - 1) + 3n; n > 1

OpenStudy (anonymous):

I was going to put C

OpenStudy (anonymous):

C. is correct

OpenStudy (anonymous):

ty :)

OpenStudy (anonymous):

yw.